# Joint-GCG: Unified Gradient-Based Poisoning Attacks on Retrieval-Augmented Generation Systems

### v1

@@ A.1 Motivation & Overview
A.2 Autoencoder Architecture & Training
@@
with . We use AdamW with cosine annealing (initial learning rate ) for 500 epochs at most, with an early stopping on validation loss.
A.3 Transfer Matrix via Least Squares
After autoencoder training, we compute a linear projection matrix to map retriever token gradients into the generator’s gradients space as below:
1. Encode Generator Embeddings: We first compute the encoded embedding for the generator embedding to match embedding dimensions and :
2. Solve Least Squares: We employ the least squares method to find the optimal projection matrix between tokens. Specifically, for each retriever token embedding , we find weights over generator tokens that minimize:
@@ A.4 Evaluation
Token Recall at Top-K (: We measured whether the corresponding ground truth retriever token embedding was found within the nearest neighbors in the retriever embedding space for each projected generator embedding. We report Top-1, Top-3, Top-5, and Top-10 Recall. Higher recall values signify better preservation of semantic similarity after projection.
@@
As shown in Table VII, the CVP autoencoder performs better on the validation set than all other methods. While Random Forest achieves the lowest Euclidean distance on the training set, the autoencoder generalizes better, achieving the lowest Projection Error on unseen data. Furthermore, the autoencoder significantly outperforms all Top-K Shared Token Recall baselines across all K values (1, 3, 5, and 10), reaching near-perfect recall even at Top-1 (97.95%). These results highlight the autoencoder’s effectiveness in learning a robust and semantically meaningful projection, essential for bridging the vocabulary gap between retrievers and generators.
@@
Appendix B Pseudo-Code for Gradient Tokenization Alignment (GTA)
@@
Appendix C Empirical Validation for Gradient Tokenization Alignment (GTA)
@@ C.1 Experimental Setup
C.2 Results and Discussion
The comparative results are presented in Table VIII. Across all three datasets, the full Joint-GCG framework with GTA consistently outperformed the simpler alternative, particularly in terms of and .
@@ C.2 Results and Discussion
As shown in Table VIII, while both methods achieved high , Joint-GCG with GTA demonstrated notably higher , especially on the NQ dataset (i.e., vs ). Furthermore, Joint-GCG consistently achieved a better (lower) , indicating a more effective retrieval manipulation that places the malicious document at a more prominent rank. For instance, on NQ, was for Joint-GCG compared to for the alternative method.
@@
Appendix D Effect of the Position of Poisoned Documents
Appendix E Additional Steps for Baseline
To further investigate the efficacy of our Joint-GCG framework and the baseline methods, we conducted experiments that extended the optimization steps from 64 to 128 for LIAR, aligning the optimization steps for the generator. The results of these experiments are presented in Table IX.
Analysis of Table IX in conjunction with the 64-step results (Table I) reveals that while LIAR benefits from increased steps with modest improvements in in specific scenarios (for instance, a 5% increase on NQ with Llama3 and a minor gain on MS MARCO with Qwen2), Joint-GCG remains superior. Joint-GCG achieves higher or comparable attack success rates with only 64 steps, demonstrating greater efficiency.
Appendix F Quantifying Computational Overhead
@@
| Prompt: You are a helpful assistant, below is a query from a user and some relevant contexts. Answer the question given the information in those contexts. Your answer should be short and concise. If you cannot find the answer to the question, just say "I don’t know". |

### v2

@@ B.1 Cross-Vocabulary Projection (CVP) Details
Autoencoder Architecture & Training
@@
with . We use AdamW with cosine annealing (initial learning rate ) to train the model for 500 epochs at most, with an early stopping on validation loss.
Transfer Matrix via Least Squares
After autoencoder training, we compute a linear projection matrix to map retriever token gradients into the generator’s gradient space as below:
1.
Encode Generator Embeddings: We first compute the encoded embedding for the generator embedding to match embedding dimensions and :
(19) -
2.
Solve Least Squares: We employ the least squares method to find the optimal projection matrix between tokens. Specifically, for each retriever token embedding , we find weights over generator tokens that minimize:
@@
Token Recall at Top-K (): We measured whether the corresponding ground truth retriever token embedding was found within the nearest neighbors in the retriever embedding space for each projected generator embedding. We report Top-1, Top-3, Top-5, and Top-10 Recall. Higher recall values signify better preservation of semantic similarity after projection.
@@
As shown in Table 4, the CVP autoencoder performs better on the validation set than all other methods. While Random Forest achieves the lowest Euclidean distance on the training set, the autoencoder generalizes better, achieving the lowest Projection Error on unseen data. Furthermore, the autoencoder significantly outperforms all Top-K Shared Token Recall baselines across all K values (1, 3, 5, and 10), reaching near-perfect recall even at Top-1 (97.95%). These results highlight the autoencoder’s effectiveness in learning a robust and semantically meaningful projection, essential for bridging the vocabulary gap between retrievers and generators.
@@
B.2 Gradient Tokenization Alignment (GTA)
@@
B.3 Empirical Validation for Gradient Tokenization Alignment (GTA)
@@
Results and Discussion
The comparative results are presented in Table 5. Across all three datasets, the full Joint-GCG framework with GTA consistently outperformed the simpler alternative, particularly in terms of and .
@@
As shown in Table 5, while both methods achieved high , Joint-GCG with GTA demonstrated notably higher , especially on the NQ dataset (i.e., vs ). Furthermore, Joint-GCG consistently achieved a better (lower) , indicating a more effective retrieval manipulation that places the malicious document at a more prominent rank. For instance, on NQ, was for Joint-GCG compared to for the alternative method.
@@
Appendix C Extended Experimental Details
This section provides a more granular look at the experimental setup used to evaluate Joint-GCG.
C.1 Datasets
MS MARCO (nguyen2016ms): The Microsoft Machine Reading Comprehension (MS MARCO) dataset is a large-scale benchmark for information retrieval and question answering. It consists of real-world queries sampled from Bing’s search logs, with passages extracted from web documents as candidate answers. Our experiments use some subsets of the queries determined by the baselines we compare.
Natural Questions (NQ) (kwiatkowski2019natural): This dataset consists of naturally occurring questions posed by users in Google Search, with human-annotated answers extracted from Wikipedia articles. Unlike MS MARCO, which focuses on search engine queries, NQ emphasizes the extraction of factual knowledge from structured sources. The dataset is particularly useful for evaluating a system’s ability to retrieve and extract concise answers from a large-scale knowledge base.
HotpotQA (yang2018hotpotqa): A multi-hop question-answering dataset that requires reasoning over multiple documents to arrive at a correct answer. Unlike single-hop QA datasets, where the answer is typically found within a single passage, HotpotQA demands integrating information across different documents, making it an excellent benchmark for assessing the model’s capability to handle complex reasoning tasks.
The dataset used for comparison with PoisonedRAG and LIAR consists of 100 queries, each sampled by PoisonedRAG from the MS MARCO, NQ, and HotpotQA datasets.
For comparison with Phantom, we used the same dataset of 25 queries for each trigger. Each query was sampled by Phantom from the MS MARCO dataset and was selected to contain a particular trigger word.
C.2 Retriever Models
Contriever (izacard2021unsupervised): A contrastive learning based retrieval model designed for unsupervised sentence embeddings. Contriever is trained using a contrastive loss, which encourages similar text pairs to have closer embeddings while pushing apart unrelated pairs. This approach has demonstrated strong performance in retrieval tasks, particularly in zero-shot and low-resource settings, making it a robust choice for open-domain QA scenarios. We utilize the contriever-msmarco model here.
BGE (BAAI General Embedding) (xiao2024c): A family of embedding models developed by the Beijing Academy of Artificial Intelligence (BAAI), optimized explicitly for retrieval tasks. BGE models are trained using large-scale datasets designed to produce highly efficient representations, facilitating fast and accurate retrieval in Approximate Nearest Neighbor (ANN) search. Their effectiveness in dense retrieval tasks makes them a competitive alternative to traditional retrieval methods. We utilize the BGE-base-en-v1.5 model here.
C.3 Generator Models
Llama3-8B (dubey2024llama): A cutting-edge open-source LLM developed by Meta AI featuring 8 billion parameters. As a successor to the highly successful Llama and Llama 2 models, Llama3 is designed to offer superior reasoning, comprehension, and response generation capabilities. Its accessibility and state-of-the-art performance make it a strong candidate for evaluating adversarial robustness in RAG scenarios.
Qwen2-7B (yang2024qwen2technicalreport): A 7-billion-parameter model developed by Alibaba Cloud as part of the Qwen series. Qwen2 is known for its strong multilingual capabilities and efficient inference, making it a competitive choice for real-world applications. Its training methodology emphasizes knowledge-rich responses, which makes it particularly interesting to evaluate how retrieval-augmented attacks influence factual generation and reasoning.
C.4 Metrics
Generation Attack Success Rate (): The percentage of target queries for which the LLM generates the desired target output, i.e., the generated output contains the target. We use this approach to align with PoisonedRAG (zou2024poisonedrag), as it shows negligible differences from human evaluation.
C.5 Experimental Settings
In experiments comparing Joint-GCG to PoisonedRAG with GCG and LIAR, we set the optimizable sequence length to 32 tokens and the optimization steps to 64. We employed a variant of GCG, MCG (chaudhari2024phantom), to enhance the attack efficiency and utilized its default hyperparameters. We expressly set the Batch Size to 128 and TopK to 16, used ASCII-character-only tokens for attacks, and configured the optimization target to the incorrect answer. For the LIAR method, we used a 1:1 ratio for the optimizable sequence length (16 tokens each for the retriever and generator) and optimization step count (8 steps each for the retriever and generator), attacking the retriever first, followed by the generator.
For batch query poisoning, we employed the same prescribed in Phantom for Denial-of-Service attacks and concatenated the optimizable sequence on it. Due to the time-consuming nature of the computation, we performed only one round of experiments. We set the optimizable sequence to 128 and the optimization step counts to 32. For Phantom, we followed their settings, using a retriever-optimizable sequence length of 128 and a generator-optimizable sequence length of 8. We perform 256 steps of GCG for the 128 retriever tokens in advance for both methods to ensure a robust retrieval rate on batch queries, facilitating further optimization of the generator-optimizable tokens.
To compare Joint-GCG with PoisonedRAG and LIAR, we used the same data as PoisonedRAG’s black-box retriever approach, using their first generated fake corpus111Available at PoisonedRAG’s official GitHub Repository. of towards the corresponding dataset in each experiment. We add an optimizable sequence at the beginning of the fake corpus, initialized with ”!”.
For comparisons with Phantom, the documents consist of three parts concatenated together as prescribed in their work, , , and , respectively. We first initialized with ”?” and with ”!”.
C.6 Effect of the Position of Poisoned Documents
C.7 Additional Steps for Baseline
To further investigate the efficacy of our Joint-GCG framework and the baseline methods, we conducted experiments that extended the optimization steps from 64 to 128 for LIAR, aligning the optimization steps for the generator. The results of these experiments are presented in Table 6.
Analysis of Table 6 in conjunction with the 64-step results (in main paper) reveals that while LIAR benefits from increased steps with modest improvements in in specific scenarios (for instance, a 5% increase on NQ with Llama3 and a minor gain on MS MARCO with Qwen2), Joint-GCG remains superior. Joint-GCG achieves higher or comparable attack success rates with only 64 steps, demonstrating greater efficiency.
@@ C.7 Additional Steps for Baseline
Appendix D Additional Experimental Results
D.1 Extending to Batch Query Poisoning
Table 7 showcases the results of this experiment. Joint-GCG consistently outperforms Phantom across all tested trigger keywords and optimization steps. Joint-GCG achieves significantly higher values at earlier optimization steps and plateaus at a higher success rate. For instance, on the “xbox” trigger, Joint-GCG reaches a high of 92% at step 4, whereas Phantom plateaus at a lower 80%. These results demonstrate Joint-GCG’s strong capability in batch query poisoning, achieving more effective and faster convergence to high attack success rates than Phantom. This highlights the versatility of Joint-GCG and its potential for broader applications in RAG system manipulation beyond targeted individual queries.
| Trigger | Attack / Step | 0 | 4 | 8 | 16 | 32 |
Appendix E Potential Defensive Mechanisms
The potency of Joint-GCG motivates a critical analysis of existing and potential defensive mechanisms. While a comprehensive evaluation of all possible defenses is beyond the scope of this work, we analyze Joint-GCG’s resilience against two widely-discussed and representative defensive strategies: (1) SmoothLLM (robey2023smoothllm), a state-of-the-art input perturbation defense that targets the generator, and (2) Perplexity-Based Filtering (alon2023detecting), a common content anomaly defense that targets the corpus.
Our findings indicate that defenses targeting only one stage of the RAG pipeline (either the corpus content or the generator input) are fundamentally insufficient against a unified attack like Joint-GCG. The attack’s joint optimization paradigm allows it to create poisons that are robust against these single-point-of-failure defenses.
E.1 SmoothLLM
| Retriever | Generator | Llama3 | Qwen2 | ||
| SmoothLLM | w/o | w/ | w/o | w/ | |
| Contriver | (w/o attack) | 77% | 72% | 71% | 66% |
| 94% | 53% | 96% | 56% | ||
| BGE | (w/o attack) | 87% | 85% | 85% | 82% |
| 87% | 47% | 92% | 41% | ||
As demonstrated in Table 8, while SmoothLLM reduces Joint-GCG’s , the attack remains alarmingly potent even under this defense. For Contriever-retriever systems, Joint-GCG achieves 53% on Llama3 and 56% on Qwen2 with SmoothLLM enabled.
The persistent success of the attack under defenses highlights Joint-GCG’s robustness. Even when attenuated by SmoothLLM, the attack success rates remain comparable to undefended performance of prior methods (e.g., GCG’s around 70% in main paper). This result highlights a fundamental weakness: defenses that target only the generation stage are insufficient, as they ignore the retrieval-manipulation aspect of the attack, underscoring the challenges for defensive mechanisms and necessitating new research into retrieval-aware adversarial filtering.
E.2 Perplexity-Based Filtering
Perplexity-based filtering involves using the perplexity score to assess the likelihood of the corpus and filtering out documents that exceed a certain threshold. This approach aims to reduce the noise introduced by less relevant or spurious information retrieved by the system. To test this, we re-ran the attack with an adaptive constraint during the GCG optimization loop. We modified the greedy selection to only accept token substitutions that keep the poisoned document’s perplexity below a threshold (set to the 95th percentile of the benign corpus). This simulates an attacker who is aware of the perplexity defense and actively tries to evade it. As shown in Figure 10, the perplexity distribution of Joint-GCG optimized adversarial examples is noticeably shifted towards higher perplexity values than the actual MS MARCO corpus. This suggests that perplexity could be a potential indicator for identifying and filtering out adversarial examples.
| Settings | ||
| Joint-GCG w/ PPL Constraint | 100.00% | 73.33% |
| w/o optimize | 100.00% | 49.00% |
Table 9 presents the attack success rates. Remarkably, even when optimized with a perplexity constraint, Joint-GCG maintains a significantly higher than the baseline scenario where no optimization is performed, with a 73.33% for Joint-GCG with the perplexity constraint, compared to only 54.00% without optimization. This apparent increase demonstrates that perplexity-based filtering, in its simplest form, is inadequate against Joint-GCG. Even when adversarial examples are crafted to have perplexity values within the normal range, the attack remains potent and surpasses the baseline ASR by a large margin. This highlights the need for more advanced defenses to detect adversarial examples beyond simple perplexity thresholds.
E.3 Discussion on Advanced and Adaptive Defenses
The evaluations above, while representative, point to a larger conceptual challenge for RAG security. We briefly discuss the implications for more advanced defenses and adaptive threat models.
On Retrieval-Aware Defenses
Defenses such as paraphrase-based document sanitization upon ingestion or retrieval-time anomaly detection (e.g., checking for unusual query-document similarity) are conceptually more robust than the ones we tested. However, they are not a panacea. The Joint-GCG framework could theoretically be extended to model such defenses. For instance, an adaptive attacker could incorporate a differentiable approximation of a paraphrasing function directly into the optimization loop, explicitly finding poisons that maintain their malicious properties post-paraphrasing.
On Detection Feasibility and Threat Model
Our threat model assumes an attacker can inject documents, a prerequisite for any corpus poisoning attack. Our work focuses on maximizing the efficacy and stealth of that injected content. Orthogonal, systems-level defenses are therefore crucial for addressing the injection vector itself:
Corpus Governance: Strict access control, content-source verification (e.g., digital signatures), and immutable versioning are paramount for auditable and trusted knowledge bases.
Behavioral Monitoring: An attacker’s optimization process, particularly one leveraging corpus observations (as in our synthetic corpus experiment), may require numerous probing queries. Detecting such anomalous query patterns or sudden high-volume contributions from a new source could serve as a valuable detection signal.
Deployment Recommendations
Our research strongly suggests that no single-layer defense will be sufficient. A robust, defense-in-depth strategy is necessary, addressing each stage of the RAG pipeline:
1.
Prevent Injection (System-Level): Implement strict corpus governance and access control.
2.
Filter Content (Retrieval-Level): Employ retrieval-aware defenses such as content sanitization, paraphrase detection, or ensemble retrieval from multiple, sandboxed corpora.
3.
Harden Generation (Generator-Level): Utilize input perturbation (like SmoothLLM) and monitor outputs for policy violations or unexpected behaviors.
Joint-GCG demonstrates that a failure in any one of these layers can be decisively exploited by a sophisticated attacker who understands and optimizes against the entire RAG pipeline, reinforcing the need for a holistic security posture.
@@ E.3 Discussion on Advanced and Adaptive Defenses
| Prompt: You are a helpful assistant, below is a query from a user and some relevant contexts. Answer the question given the information in those contexts. Your answer should be short and concise. If you cannot find the answer to the question, just say ”I don’t know”. |
@@
Appendix F Limitations and Future Works
F.1 Computational Overhead
F.2 Cross-Domain Generalization
While we demonstrate strong performance across multiple QA datasets, the generalization of Joint-GCG to other domains (e.g., code generation (rani2024augmenting), medical applications (ye2024exploring), and tool-calling agents (wang2025alliesadversariesmanipulatingllm)) and different types of RAG architectures requires further investigation. Future work could explore domain-specific adaptations of the framework and evaluate its effectiveness across a broader range of applications and model architectures.
Appendix G Example of a Successful Attack
See Table 10.